A cell-free extract of a lactobacillus bacterium t
Answer #1

Glucose + 2 NAD^{+} + 2 ADP \rightarrow^{x_{1}} 2 Pyruvate + 2H_{2}O +2NADH + 2ATP

2 Pyruvate + 2 NADH \rightleftharpoons ^{x_{2}}_{x_{3}} 2 Lactate + 2 NAD^{+}

2ADP + 2P_{i} \rightarrow^{x_{4}} 2 ATP

The above three equations summerise the given system. Where x1, x2,x3, x4 are rate constants of the given equations

Also, as Lactate is periodically being removed it cannot be converted back and it can be safely assumed that x3 to be zero. Also it is clear that the product formation is non-growth associated.

Material balances can be written as

d[NADH]/dt = x1 - x2 -x3 = 0 --(1) As NADH remains constant in the system with out being generated or consumed.

d[Lactate]/dt = x2 - x3 = x2 --(2) as Lactate is continuosly being removed from the system.

Substituting (2) in (1) d[NADH]/dt = d[Lactate]/dt --(3)

From first Chemical reaction d[NADH]/dt = 2 d[Glucose]/dt --(4)

Therefore the amount of Lactate formed twice the rate of consumption of Glucose.

Assuming ideal conditions when all glucose is converted to Lactate, the amount of Lactate formed is maximum.

==> 200mM Glucose = 400mM Lactate.

In a general case the amount of glucose consumed is dependent on the rate constants of the reactions. As they are not mentioned in the problem they are idealized.